### v1

# Advancing Human-Machine Teaming: Concepts, Challenges, and Applications

Abstract.
Human-Machine Teaming (HMT) is revolutionizing collaboration across domains such as defense, healthcare, and autonomous systems by integrating AI-driven decision-making, trust calibration, and adaptive teaming. This survey presents a comprehensive taxonomy of HMT, analyzing theoretical models, including reinforcement learning, instance-based learning, and interdependence theory, alongside interdisciplinary methodologies. Unlike prior reviews, we examine team cognition, ethical AI, multi-modal interactions, and real-world evaluation frameworks. Key challenges include explain-ability, role allocation, and scalable benchmarking. We propose future research in cross-domain adaptation, trust-aware AI, and standardized testbeds. By bridging computational and social sciences, this work lays a foundation for resilient, ethical, and scalable HMT systems.
Human-machine teaming, human-AI collaboration, human-autonomy collaboration
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Collaborative interaction
1. Introduction
Human-Machine Teaming (HMT) is transforming technology, enabling seamless collaboration between humans and intelligent systems in defense, healthcare, autonomous systems, and industrial automation (Christopher Brill et al., 2018; Henry et al., 2022). For example, in autonomous military operations, HMT enables AI-driven decision support for surveillance, target recognition, and real-time strategy adjustments while maintaining human oversight. In healthcare, AI-assisted diagnostic systems and robotic surgical assistants enhance precision and efficiency, improving patient outcomes through adaptive human-AI collaboration (Gillespie, 2022). As AI advances, machines must act as effective teammates rather than mere tools (Greenberg and Marble, 2023). Despite its potential, HMT research faces challenges, including trust calibration, decision-making autonomy, adaptability, and ethical considerations (Dreslin and Baweja, 2023). Existing studies focus on specific HMT aspects like multi-robot control, human-agent trust, and cognitive load, but a comprehensive interdisciplinary understanding remains fragmented (Vats et al., 2024; Patel et al., 2024).
Primary goal of this survey. This work aims to provide a structured and holistic analysis of HMT research by integrating insights from both computational and social sciences. Unlike prior reviews that focus on individual components, such as metrics, safety, or specific teaming models, this work systematically categorizes HMT concepts, challenges, and applications (Damacharla et al., 2018; Cleland-Huang et al., 2022). This survey establishes a taxonomy of HMT systems, analyzes empirical studies on team performance, and examines key theoretical frameworks that drive human-machine collaboration. Furthermore, it presents an evaluation framework encompassing testbeds, performance metrics, and datasets, addressing the need for standardized benchmarking in HMT research (Damacharla et al., 2018).
Scope of this survey. This work primarily emphasizes trust, decision-making, cognitive alignment, and evaluation methodologies in human-machine collaboration. The survey focuses on HMT applications in defense, healthcare, robotics, and human-autonomy teaming systems while not covering purely technological advancements in AI algorithms or robotics hardware, as these aspects are extensively reviewed in engineering literature. Instead, this work highlights the interdisciplinary and human-centered dimensions of HMT to guide future research directions and foster interdisciplinary collaboration. The insights provided in this survey will help researchers, engineers, and policymakers design more effective, resilient, and trustworthy HMT systems for real-world applications.
1.1. Comparison with Existing Surveys
Table 1. Comparison of our Survey Paper with the Existing Surveys of Human-Machine Teaming Networks
Key Criteria (Murphy and Schreckenghost, 2013) (2013) (Chen and Barnes, 2014) (2014) (Lasota et al., 2017) (2017) (Damacharla et al., 2018) (2018) (Gomez et al., 2019) (2019) (Warren and Hillas, 2020) (2020) (Kashima et al., 2022) (2022) (Yang et al., 2022) (2022) (Gebru et al., 2022) (2022) (O’Neill et al., 2022) (2022) (Berretta et al., 2023) (2023) (Vats et al., 2024) (2024) Ours
Key concepts
Key performance factors
Relationships of the key components
HMT architecture
Trustworthiness of an HMT system
Extensive review of diverse HMT research and its analysis
Key theories for HMT research
Concepts, challenges, and applications of an HMT system
Metrics for HMT research
Testbeds for HMT research
Limitations/Insights
Future Research Directions
✔: Fully addressed; ▲: Partially addressed; ✘: Not addressed at all.
Murphy and Schreckenghost (2013) reviewed 42 human-robot interaction (HRI) metrics across human, robot, and system categories, identifying gaps in measuring specialized features that limit generalizability and scalability. Conducted in 2013, their survey excludes recent HRI metrics and focuses on agents’ cognitive and autonomous capabilities. Chen and Barnes (2014) surveyed human performance issues in multi-robot control, emphasizing user-interface design for enhancing human-agent teaming (HAT). However, their findings, derived from controlled lab settings, may not fully generalize complex real-world scenarios. Lasota et al. (2017) examined safety and robustness strategies in HRI, highlighting the impact of unpredictable environments, human dynamics, and sensing limitations on system states. While technical safety was the focus, the study largely omitted psychological and social aspects like user experience and trust, which are salient or influential elements when humans are involved.
Damacharla et al. (2018) categorized HMT system metrics but did not address evaluation methods, testbeds, or broader applications. Gomez et al. (2019) reviewed approaches for optimizing cyber task performance, covering automatic platform assessments, shared mental models, provenance tracking, and user-performance modeling, offering a research roadmap but predating recent developments. Warren and Hillas (2020) analyzed Lethal Autonomous Weapons Systems (LAWS) and HMT in military contexts, focusing on trust, reliability, and AI regulation but lacking empirical validation. Kashima et al. (2022) examined human-AI trust via Reliability, Availability, and Serviceability (RAS) and ethical considerations but primarily addressed one-way trust in AI, overlooking bidirectional trust and AI’s ethical obligations.
Yang et al. (2022) provided a comprehensive review of perception, decision-making, and execution approaches in HMC for robotics, emphasizing safety, efficiency, and collaboration in remote operations, surgery, and assembly. However, evaluation methodologies, such as testbeds and simulations, were not covered. Gebru et al. (2022) examined trust measurement in humans (e.g., heart-rate variability and brain activity) and machines (e.g., robustness, explainability, verifiability) but did not explore broader system resilience. O’Neill et al. (2022) reviewed 76 human-autonomy teaming studies, analyzing research settings, variables, and team composition but found most studies were lab-based with limited real-world applicability. Their survey lacked resilience, safety metrics, and cybersecurity concerns. Berretta et al. (2023) identified five HAT research clusters, i.e., human factors, task variables, AI explainability, AI-driven robotics, and AI’s impact on human perception, but omitted technology-centered perspectives on HMT advancements. Vats et al. (2024) explored Large Pre-trained Models in HAT, focusing on AI model improvements, ethical considerations, and applications, but did not cover broader HMT studies.
Table 1 compares existing HMT surveys based on key criteria. Murphy and Schreckenghost (2013) covered metrics but omitted architecture, resilience, and safety. Chen and Barnes (2014) partially addressed trust and applications but lacked testbeds, while Lasota et al. (2017) focused on safety without discussing cybersecurity. Damacharla et al. (2018) and Gomez et al. (2019) offered limited insights into resilience and cybersecurity. Warren and Hillas (2020) and Kashima et al. (2022) prioritized trust over technical evaluations. Yang et al. (2022) covered architecture but omitted trust and resilience, whereas Gebru et al. (2022) emphasized trust measurement over system resilience and safety. O’Neill et al. (2022) analyzed human-autonomy teaming but overlooked resilience and cybersecurity. Berretta et al. (2023) examined human-AI teaming from a human-centric perspective, while Vats et al. (2024) focused solely on Large Pre-trained Models in HAT.
Our review fully addresses these dimensions, covering architecture, trust, resilience, safety, metrics, testbeds, multi-domain applicability, and cybersecurity, providing a unified framework. We offer in-depth architectural insights, robust evaluation methods, and emerging threat analyses for HMT systems, delivering actionable recommendations for real-world implementations and future research.
1.2. Key Contributions
Our survey paper made the following key contributions:
Comprehensive Taxonomy. This survey establishes a structured taxonomy of Human-Machine Teaming (HMT), categorizing key components, architectures, and trustworthiness factors. Unlike prior reviews focused on specific HMT subdomains or applications, our taxonomy provides a unified framework integrating diverse HMT systems, including human-unmanned vehicle teaming, human-robot collaboration, and AI-driven social media interactions. This broader, systematic classification enables researchers to uncover cross-disciplinary insights that were previously overlooked.
Empirical Study Insights. We analyze empirical studies across multiple HMT dimensions, including team training, autonomy, trust, and shared mental models (SMMs). Compared to existing reviews that primarily focus on conceptual discussions or theoretical models, our work synthesizes findings from both experimental and real-world deployments. By identifying critical success factors such as communication effectiveness, situational awareness, and ethical considerations, this review highlights research gaps and calls for standardized methodologies to improve reproducibility across studies.
Evaluation Framework. This survey presents an in-depth discussion of HMT evaluation methodologies, covering testbeds, performance metrics, and dataset availability. Unlike prior reviews emphasizing qualitative or quantitative assessment techniques, we integrate both perspectives, proposing a holistic evaluation framework. Our review uniquely emphasizes the role of trust, resilience, and ethical constraints in performance assessments, ensuring a more robust and context-aware evaluation of HMT implementations.
Multidisciplinary Perspective. Unlike many prior surveys focusing predominantly on computational or human-centered aspects, our review bridges both perspectives by incorporating social science theories and AI-driven frameworks. We explore how Human-in-the-Loop Reinforcement Learning (HRL) and Instance-Based Learning Theory (IBLT) can enhance adaptive teaming mechanisms. By synthesizing insights from cognitive science, machine learning, and human factors research, this survey offers a more integrative approach to fostering synergistic human-AI collaboration.
Figure 1. Structure of this review paper.
Comparative Strength over Prior Surveys. Unlike existing HMT reviews, such as (Murphy and Schreckenghost, 2013) on metrics, (Chen and Barnes, 2014) on multi-robot control, and (Lasota et al., 2017) on safety, our work provides a broader, more inclusive analysis covering architecture, performance metrics, trust dynamics, and interdisciplinary evaluation. While surveys like (Warren and Hillas, 2020; Kashima et al., 2022) explore trust, they do not fully examine its relationship with team coordination and task allocation in HMT. Our review fills these gaps by systematically comparing diverse HMT applications and integrating key concepts into a unified research framework.
Future Research Roadmap. We outline key future research directions, including enhancing explainability in HMT systems, improving AI-human trust calibration, and mitigating ethical and bias-related challenges. Unlike previous reviews that primarily summarize past findings, our work proposes concrete steps for advancing HMT research. We emphasize the need for cross-sector collaboration between academia, industry, and government to develop standardized benchmarks and evaluation protocols. This roadmap serves as a guideline for ensuring the practical deployment of HMT across various domains, including defense, healthcare, and industrial automation.
The remainder of this paper follows the structure outlined in Fig. 1.
2. Taxonomies of HMT Systems
2.1. Related Concepts and Key Factors in HMT
Concept of HMT. HMT has been discussed across various related domains, with their definitions discussed below.
  • Human-Machine Interaction (HMI). The interaction and communication between humans and machines (e.g., industrial or transportation systems) via an interface (Johannsen, 2009; Hoc, 2000), where the human can act as a decision-maker.
  • Human-Machine Collaboration (HMC). A collaborative form of HMI where humans and machines work toward shared goals through interaction, common in health management, surveillance, and manufacturing (Pizoń and Gola, 2023).
  • Human-Machine Teaming (HMT). The partnership of humans and machines collaborating to achieve a shared objective through interaction and communication (Lyn Paul et al., 2019).
  • Human-Machine Teaming Network (HMTN). A network of HMTs collaborating under a structured framework (e.g., Cloud-Sharing Network) (Longa et al., 2022), with each HMT serving as a specialized team member rather than a simple swarm.
  • Human-Machine Symbiosis (HMS). A mutually beneficial relationship between humans, offering subjective expertise, and machines, providing objective computational abilities (Gill, 2012). This collaboration unites complementary strengths to support complex decision-making in human-centered systems.
  • Human-Autonomy Teaming (HAT). The interdependent collaboration of humans and autonomous agents, with distinct roles, working toward shared goals (O’Neill et al., 2022). HAT agents independently interact with other team members.
  • Human-AI Teaming Systems (HATS). An integrated system of humans and AI models collaborating to achieve shared objectives by combining their respective strengths (Andrews et al., 2023).
  • Human-Robot Systems (HRS). The physical collaboration between humans and intelligent devices, such as sensors and actuators, combining human dexterity and robotic precision for improved productivity (Ogenyi et al., 2021).
  • Autonomous Systems. Systems capable of adapting to unforeseen events during operation (Watson and Scheidt, 2005), featuring self-perception, self-planning, and self-learning. A joint unmanned air combat system exemplifies these capabilities (Chen et al., 2021).
  • Automation & Autonomy. Autonomy, a higher level of automation, is defined by self-determination and independent operation without human intervention (Chen et al., 2021), unlike automation, which depends on external control.
Hierarchical Relationships in HMT Systems.
Figure 2. Hierarchical diagram illustrating relationships and dependencies among HMT-related concepts.
Fig. 2 illustrates the interconnections among key HMT concepts. At the center, HMT represents collaborative human-machine interactions for shared objectives. Surrounding it, HMI, HMC, HAT, HMS, HATS, and HRS contribute to different collaboration aspects. Broader concepts like HMTN, AS, and AA provide higher-level structures. The diagram illustrates the hierarchical relationships, providing a clearer understanding of their roles in HMT.
Comparative Analysis of HMT Concepts. Table 2 compares key HMT concepts based on collaboration level, autonomy, human involvement, and system focus. Each concept is briefly described, with corresponding attributes in other columns. The table spans basic HMI to advanced human-autonomy and HAT systems (HATS), illustrating varying levels of collaboration, autonomy, and human involvement. The system focus column identifies primary applications, such as decision-making, autonomous systems, or robotics.
Table 2. Comparison of Human-Machine Teaming Concepts
Concept Description Collaboration Level Autonomy Human Involvement
Human-Machine Interaction (HMI) Interaction between human and machine via interface. Low Low Decision-making
Human-Machine Collaboration (HMC) Humans and machines working together for shared goals. Medium Medium Active
Human-Machine Teaming (HMT) Partnership between humans and machines working toward a shared objective. High Low to Medium High
HMT Network (HMTN) Network of HMTs collaborating under a framework. High Low to Medium High
Human-Machine Symbiosis (HMS) Mutually beneficial collaboration leveraging human expertise and machine precision. High Low High
Human-Autonomy Teaming (HAT) Collaboration between humans and autonomous agents. High High High
Human-AI Teaming Systems (HATS) Integrated system of humans and AI models working toward shared objectives. High High High
Human-Robot Systems (HRS) Physical collaboration between humans and robots. Medium Medium Medium
Autonomous Systems (AS) Systems that operate independently and adapt to events. Low to Medium High Low
Automation & Autonomy (AA) Automation is external control, while autonomy is a self-determined operation. Low Medium to High Low
Key insights from Table 2 include the increasing complexity of systems as collaboration, autonomy, and human involvement grow. While HMI centers on human control, HMC, HMT, and HAT introduce progressively higher collaboration and autonomy, with HAT and HATS integrating human and autonomous agents. In addition, Table 2 highlights how these concepts apply across domains, from human-robot collaboration in HRS to decision support in HMS, showcasing the broad scope of human-machine systems.
Key Factors for Enhancing HMT Systems. We identify the following for enhancing HMT systems’ performance:
Leadership: In HMTS, human team leaders must communicate effectively to manage diverse teams (Flathmann et al., 2021). Strong communication aids information search, analysis, and resource management, ensuring tasks are completed on time.
Team Diversity: Trustworthy AI systems require diverse teams unified by shared ethics (Smith, 2019). Diversity in gender, race, education, skills, and problem-solving reduces bias and unintended consequences, leading to fairer solutions.
Figure 3. Conceptual relationships of the key factors influencing the performance of HMT systems.
Shared Mental Model (SMM): An SMM represents a team’s shared understanding of system behavior, enhancing coordination (Andrews et al., 2023). It consists of (1) the task model, which defines goals and processes, and (2) the team model, which outlines roles and expectations. SMM benefits include minimizing miscommunication, fostering trust, and improving coordination.
Trust: Trust in HMTS is studied in: (1) Trust in automation, an agent’s reliability in uncertain conditions, assessed through trust metrics (Jian et al., 2000). (2) Mutual trust, which enhances communication, consistency, shared knowledge, and adaptability for efficient teamwork.
Fig. 3 represents a conceptual relationship map of key factors influencing the performance of HMT systems. It illustrates the interactions between four main concepts: Leadership (LS), Team Diversity (TD), Shared Mental Model (SMM), and Trust. Leadership (LS) influences both SMM and Trust, promoting team alignment and transparency. Team Diversity (TD) enhances SMM by integrating diverse perspectives while increasing trust by fostering mutual respect among team members. A well-established SMM improves Trust by reducing misunderstandings and increasing predictability in interactions. The arrows labeled enhance show that Leadership and Diversity contribute to a stronger Shared Mental Model, while the arrows labeled increase indicate that Leadership and Diversity also increase Trust. This map highlights the interdependencies of these factors and their combined effect on HMT system performance.
2.2. Key Components of HMT Systems
To develop an HMT system, recent studies have identified the following core components:
  • Agents: Members of a team consisting of humans and machines where (1) human agents are categorized as operators, engineers, maintenance personnel, or managers, responsible for controlling machines or making decisions (Johannsen, 2009), and (2) machine agents are also called a bot or AI agent, refers to any dynamic technical system (e.g., autonomous systems, decision-making, or decision support systems) (Johannsen, 2009).
  • Role: A container for an agent with the required qualifications (Madni and Madni, 2018), assigned based on task-specific requirements.
  • Inputs: Categorized by their focus, such as humans, machines, or contexts (Stowers et al., 2017). (1) Human input: Includes cognitive competencies (e.g., skills, memory, expertise) and interpersonal traits (e.g., personality, values), impacting interaction quality. (2) Machine input: Includes automation level, reliability, and transparency, influencing trust, workload, and system performance. (3) Contextual input: Covers environmental factors and task variables, including external conditions, task type, complexity, and load.
  • Human-Machine Interface (HMI): Enables interaction between humans and cyber-physical elements through displays, controls, and information exchange. Effective HMIs support situational awareness, transparent machine operations, appropriate intervention, and manageable cognitive workload (Madni and Madni, 2018; Wachs and Duerstock, 2010).
  • Task: A job performed by an agent assigned to a role. Humans and machines share task loads and switch roles as needed. Tasks are reallocated based on outcomes and evaluations in dynamic environments (Madni and Madni, 2018; Hu and Chen, 2017).
  • Environment: Contains factors and conditions affecting task performance. In dynamic environments, task allocation depends on context, availability of agents, and human cognitive and emotional states (Madni and Madni, 2018).
  • Disrupting Events: External or internal disturbances requiring human-machine teams to adapt. Such events can impact environments, altering consequences and task sharing. Models must account for these disruptions to manage task allocation effectively (Madni and Madni, 2018; Connors, 1998).
  • Adapting Actions: Modifications to structure, sequence, or HMT parameters in response to disruptions. Effective adaptation prevents poor outcomes and improves task evaluations, ensuring robust performance (Madni and Madni, 2018; Dubois and Le Ny, 2020).
2.3. Architecture of an HMT System
Queuing Network-Model Human Processor (QN-MHP). This is designed to model multitask human performance by integrating queuing networks with symbolic cognitive models. It effectively models dual tasks, such as in-vehicle map reading and steering, demonstrating its potential for cognitive modeling. Despite its novel approach, QN-MHP lacks accuracy in predicting task durations under certain conditions, such as varying glance durations on sharp curves, and does not address speed control or complex road geometry adjustments (Liu et al., 2006).
Cyber-Physical Human System (CPHS). This explored options in simulated contexts. The architecture integrates human electrophysiological signals to infer intent and sensors to capture environmental changes processed by a controller. This feedback loop mirrors an adaptive CPHS with tight coupling and strict temporal constraints (Madni and Madni, 2018).
HMI for Active and Assisted Living Solutions. This approach emphasizes HMI as a critical factor in adopting active and assisted living solutions. Using a context-based approach, it aims to enhance HMS efficacy, adaptability, and performance. Experiments with the INRIA dataset demonstrated that adaptable actions to dynamic contexts improve system performance when algorithms are sequenced. However, the approach is impractical for complex systems requiring expert analysis and lacks details on distributed programming for large-scale systems (Quintas et al., 2017).
Assistive Robotic System for People with Physical Disabilities. This system enables simple tasks through a semi-autonomous decision-and-control architecture. It features soft robotics, combinable skill libraries, assistive sensing, and multimodal feedback. Components include lightweight design, impedance control, collision detection, and virtual workspace boundaries. However, it is limited to simple tasks and cannot handle complex scenarios (Vogel et al., 2015).
PREdictive SENsorimotor Control and Emulation Architecture (PRESENCE). The PRESENCE architecture was introduced for natural speech interactions in HMI. Inspired by neurobiology, it integrates hierarchical feedback control based on perceptual control theory (PCT), where speakers and listeners align perceptions and intentions through closed-loop feedback. While innovative at the time, the study’s settings are outdated compared to modern HMI research (Moore, 2007).
Figure 4. Architecture of a human-machine teaming system based on its core components.
Proposed HMT Architecture. Building on the core components of HMT systems, we designed an HMT architecture, illustrated in Fig. 4, where each component is essential for optimizing task and team performance. This architecture integrates human and machine agents, an HMI for bidirectional communication, and SMMs to align team roles and task objectives. Inputs from human, machine, and contextual sources and continuous feedback loops enable dynamic adaptability to environmental and task changes. The architecture emphasizes trustworthy interactions, leveraging the synergy of human cognitive capabilities and machine-driven precision to achieve effective task execution. Fig. 4 demonstrates a comprehensive framework for enhancing HMT systems’ operational efficiency and resilience by addressing adaptability and trust as central elements.
2.4. Key Components of Trustworthy HMT Systems
Developing trustworthy HMT systems requires addressing key challenges (Lyn Paul et al., 2019). Understanding user needs in dynamic contexts is essential, particularly for complex tasks like cybersecurity. Seamless collaboration depends on integrating human decision-making models with machine analytics. AI security and reliability must be ensured by mitigating adversarial attacks and biases. Transparency and explainability foster trust between human and machine agents. Successful collaboration in high-pressure environments depends on strong interactions and a shared team framework to improve communication and coordination. Building on these challenges, we discuss the key factors necessary for developing trustworthy HMT systems, focusing on their role in enhancing reliability and collaboration.
Key Properties of Trustworthy HMT Systems. Via extensive literature review, we identified the following properties essential for building a trustworthy HMT system:
  • Team Resilience: Resilience ensures robustness in industrial production and preparedness for disruptions. Adaptability is key, requiring smooth adjustments in responsibilities and work divisions. Effective HMT systems must support flexible task allocations and role changes while remaining resilient to environmental and system changes (Kaasinen et al., 2022).
  • Team Fairness: Fairness in HMT systems emphasizes non-discrimination and unbiased AI models that are inclusive and representative (Pflanzer et al., 2023). Fairness metrics for human-robot teams include equality of workload (equal subtasks regardless of type or capability), equality of capability (matching subtasks by time, accuracy, and difficulty), and equality of task type (equal distribution of subtasks within the same category) (Chang et al., 2020).
  • Team Proactivity: Machine behavior can be reactive, active, or proactive (Edgar et al., 2023). Reactive agents respond to requests, active agents provide information as needed, and proactive agents initiate tasks, remind teammates, and adapt to situations. Proactive robots improve efficiency and performance in demanding environments (Edgar et al., 2023).
  • Team Usability: Team usability refers to how effectively humans collaborate with AI to achieve goals (Smith, 2019). High usability fosters trust and willingness to engage with AI teammates, enhancing human-machine collaboration.
  • Team Trust: Team trust reflects humans’ confidence in autonomous systems and willingness to interact (Ibrahim et al., 2022). Trust is influenced by system reliability, transparency, predictability, and proactive assistance (Lyons et al., 2019). Mutual trust requires machines to estimate human trust levels and adapt behavior to ensure safety and interaction quality (Alhaji et al., 2020).
  • Team Accountability: Accountability entails shared responsibility between humans and AI, ensuring transparency and enabling evaluation of decisions and actions (Chen and Barnes, 2014; Lopez et al., 2023).
  • Team Security: Team security is the ability of human and machine teammates to counter attacks while maintaining performance. Key aspects include machine involvement, role transparency, and action traceability (Assaad and Boshuijzen-van Burken, 2023; Brown, 2019). Simple decision rules and high machine involvement enhance security, but poor coordination weakens it.
  • Team Agility: Agility measures a team’s ability to adapt to changing task requirements and unforeseen obstacles (Demir et al., 2021a). Agile teams emphasize collaboration over individual assignments, enabling interactive decision-making and adaptability without requiring a designated leader.
According to the metric framework STRAM (Cho et al., 2019), which defines the key metric attributes of a computer-based trustworthy system (Security, Trust, Resilience, and Agility), we define the trustworthiness of an HMT system based on team security, trust, resilience, and agility, as shown in Fig. 5. These attributes form the backbone of a reliable and effective human-machine teaming framework. Meanwhile, auxiliary attributes like team usability, fairness, accountability, and proactivity complement and enhance these core elements through specific interdependencies.
Team usability enhances agility by ensuring intuitive human-machine interactions that adapt quickly to changing tasks while supporting resilience by reducing cognitive load and improving coordination under stress. Fairness bolsters trust and resilience through equitable task allocation and unbiased decision-making, preventing overload and maintaining balance.
Figure 5. Attributes of a trustworthy HMT system based on the STRAM framework.
Accountability reinforces trust and security by enabling transparent, traceable decisions, mitigating errors, and enhancing reliability. Proactivity drives agility and resilience by dynamically anticipating changes and addressing challenges, fostering trust through foresight and reliability.
These auxiliary attributes collectively strengthen the core dimensions of team security, trust, resilience, and agility. For example, usability and proactivity enhance agility, while fairness and accountability reinforce trust and resilience. Recognizing these interdependencies is key to designing HMT systems that combine technical robustness with ethical integrity, enabling effective human-machine collaboration in complex environments.
3. Empirical Studies to Promote Team Performance
This section categorizes subsections by key HMT topics that enhance team performance. Detailed discussions of reviewed papers are in the supplement (Appendix A) due to space constraints.
3.1. Team Training
Key Ideas. Johnson et al. (2021) examined the impact of communication and calibration on HAT performance using a mixed testbed of real and simulated platforms. They tested three training approaches, including control, coordination, and calibration, showing that coordination improved trust but did not outperform control-trained teams. Limited statistical power constrained conclusions. Harpstead et al. (2023) applied speculative game design to study HMT dynamics, focusing on roles, structures, and coordination. While effective for exploring team interactions, the approach lacked details on AI training and relied on simplified game environments.
Myers et al. (2019) introduced autonomous synthetic teammates (ASTs) to reduce scheduling and training costs, balancing human-like behavior with task performance. However, cognitive fidelity limitations affected their effectiveness in complex tasks. O’Neill et al. (2023) proposed the Input-Mediator-Output (IMO) model to structure HMT research, though it overlooked bidirectional feedback loops essential in dynamic interactions. Endsley and Appleby (2024) developed an HMT taxonomy advocating for collaborative autonomous systems over autonomy-centric designs. Cleland-Huang et al. (2024) extended the MAPE-K feedback loop to MAPE-KHMT, integrating adaptive human-machine interactions. Lematta et al. (2024) highlighted HMT assurance challenges, using General Motors Cruise robotaxis to illustrate trust and safety failures due to poor design, impacting emergency operations and human interaction reliability.
Merits. The existing approaches to team training using HMT have advanced HMT across multiple domains, focusing on collaboration, communication, and decision-making. A key theme is structured communication and calibration, where training enhances situational awareness and trust in AI teammates (Johnson et al., 2021). Similarly, role allocation and task adaptability have been explored in speculative game design and asymmetric cooperative tasks, providing insights into team coordination and knowledge distribution (Harpstead et al., 2023). Integrating synthetic agents, such as Autonomous Synthetic Teammates (ASTs), optimizes training efficiency while balancing performance and human-like behavior (Myers et al., 2019). Methodologies for categorizing HMT interactions have been refined, such as the Input-Mediator-Output (IMO) model (O’Neill et al., 2023) and MAPE-KHMT feedback integration (Cleland-Huang et al., 2024). Advancements in HMT assurance and real-world validation are evident in AI deployment failures, such as the General Motors Cruise robotaxis case (Lematta et al., 2024). These studies highlight the importance of transparent interaction mechanisms, system reliability, and robust human-AI collaboration in safety-critical applications.
Limitations. Despite advancements, key challenges remain. Many studies rely on simulated environments like Minecraft-based experiments (Chiou et al., 2021) or grid-world tasks (Harpstead et al., 2023), limiting real-world applicability. Empirical testing in operational environments is needed to assess scalability and effectiveness. While AI transparency and trust calibration are emphasized, studies often focus on conceptual models without experimental validation. Frameworks like (O’Neill et al., 2023) and MAPE-KHMT (Cleland-Huang et al., 2024) improve HMT modeling but lack systematic deployment in real-world decision-making.
Human-centered training remains underexplored. Studies on coordination and training (Johnson et al., 2021) do not fully address non-expert adaptation, raising concerns about accessibility and usability in dynamic environments. HMT failure analysis is often case-specific without scalable solutions. (Lematta et al., 2024) highlights trust and safety failures in autonomous vehicles but lacks a structured framework for improving AI-driven systems. Addressing these challenges requires empirical validation, better AI training, and human-centered frameworks to enhance HMT systems’ robustness and viability.
3.2. Team Autonomy
Key Ideas. Recent studies examine the role of AI teammates in HMT, emphasizing performance improvement, cognitive adaptation, and optimized task execution. AI teammates have been shown to improve team coordination, mediate interactions, and positively influence human perceptions of human-AI teams (Schelble et al., 2023). Spatial awareness also plays a crucial role in team cognition, as varying levels of awareness impact mental model similarity, decision-making efficiency, and overall team performance (Schelble et al., 2022b). Beyond team cognition, adaptability is an essential trait for AI teammates, where autonomous agents must adjust their behaviors dynamically based on evolving task demands and human interaction patterns (Hauptman et al., 2023). AI-driven automation has also been leveraged in large-scale data analysis, such as in cybersecurity, where machine learning and natural language processing (NLP) techniques enable efficient taxonomy construction with reduced human effort (Mahaini et al., 2019). These studies collectively highlight the increasing integration of AI teammates in diverse domains, from emergency response coordination to cognitive adaptability and automated data classification.
Merits. The reviewed works demonstrate AI’s potential to enhance HMT performance by improving task execution, cognitive adaptability, and coordination efficiency. Studies on AI teammates reveal their ability to mediate interactions and foster human-AI collaboration, indicating that well-integrated AI systems can support human decision-making and situational awareness (Schelble et al., 2023). The examination of spatial awareness provides valuable insights into the cognitive processes underlying human-agent interactions, emphasizing the importance of shared mental models in decision-making (Schelble et al., 2022b). Further, research on autonomy levels in AI teammates highlights the need for adaptive decision-making models that align AI behavior with evolving team dynamics (Hauptman et al., 2023). Automated data analysis in cybersecurity showcases the benefits of integrating AI for large-scale information processing, demonstrating how human expertise and NLP techniques can collaboratively generate structured taxonomies with improved accuracy and efficiency (Mahaini et al., 2019). These contributions collectively advance the understanding of how AI teammates can improve human performance, decision-making, and large-scale knowledge organization in complex environments.
Limitations. Despite their contributions, the reviewed works face notable limitations. Many studies rely on controlled simulations with simplified tasks, such as abstract game-based environments or pre-defined operational conditions, which limit their applicability to real-world HMT scenarios (Schelble et al., 2023, 2022b). The reliance on experimental settings without explicit communication between AI teammates and humans also constrains the generalizability of these findings, as real-world tasks often involve more complex, bidirectional interactions. In addition, research on AI adaptability and autonomy levels is often based on hypothetical questionnaire-based studies, which may not fully capture the complexities of real-time decision-making (Hauptman et al., 2023). While AI-driven taxonomy construction in cybersecurity presents promising applications, its effectiveness in other domains remains uncertain, and its reliance on structured datasets raises concerns about adaptability to unstructured or rapidly changing information landscapes (Mahaini et al., 2019). These limitations highlight the need for empirical validation in real-world environments, better integration of communication mechanisms in human-agent teaming, and broader assessments of AI-driven automation across multiple domains.
3.3. Trust & Decision-Making
Key Ideas. Recent studies explore the role of trust, decision-making, and interaction dynamics in HMT, particularly in high-stakes environments like medical care, emergency response, and military operations. Machine learning-based systems such as the Targeted Real-time Warning System (TREWS) demonstrate effectiveness in clinical decision-making but suffer from low user trust due to a lack of explainability (Henry et al., 2022). To address such trust issues, architectures like the EASE framework (Bersani et al., 2023) integrate statistical and machine learning models to enhance transparency, while frameworks like ITA (Centeio Jorge et al., 2024) evaluate trust along competence, willingness, and external factors to optimize task allocation. Trust formation in HMT is also influenced by perception, cognitive alignment, and system reliability (Zhang et al., 2020; Demir et al., 2021b). Studies on ”swift trust” (Haring et al., 2021) suggest that appearance and functionality play crucial roles in early trust formation in human-robot teams, though long-term trust requires robust interactions. Furthermore, trust levels impact team performance, as in RPAS simulations where low trust in autonomous agents created negative feedback loops and hindered collaboration (McNeese et al., 2021a).
Merits. The reviewed works emphasize the complex interplay between trust, decision-making, and team cognition in HMT. Machine learning models such as TREWS demonstrate how AI-driven systems can enhance decision-making in critical environments (Henry et al., 2022). Studies on interaction dynamics reveal how trust evolves over time, linking positive AI experiences to increased human trust and collaboration (Zhang et al., 2020; Demir et al., 2021b). Trust assessment frameworks, including ITA and the military-oriented HAT framework (Centeio Jorge et al., 2024; Maathuis, 2024), provide structured approaches to designing reliable HMT systems by integrating ethical, societal, and operational considerations. Research on trust in team composition further highlights how mixed human-AI teams influence cognitive alignment and coordination (Musick et al., 2021; Schelble et al., 2022a), offering insights into the impact of agent presence on human cognition. These studies collectively advance the field by identifying key factors that shape trust in HMT and proposing methodologies to improve human-agent collaboration.
Limitations. Despite their contributions, these studies face notable limitations. Many rely on controlled simulations with simplified tasks, such as emergency response simulations (Schelble et al., 2022a) and RPAS scenarios (McNeese et al., 2021a), which limit real-world applicability. Small sample sizes, as seen in studies on medical AI adoption (Henry et al., 2022) and sentiment in multi-agent teams (Musick et al., 2021), reduce generalizability. Trust formation research often lacks empirical validation beyond theoretical models, as demonstrated in studies on swift trust (Haring et al., 2021). Moreover, some approaches focus on static trust measurement rather than evolving trust dynamics, overlooking how trust adapts to system learning and behavioral changes over time. The challenge of generalizing findings across domains is evident in cybersecurity-focused models like EASE (Bersani et al., 2023) and AI perception studies with unbalanced participant demographics (Zhang et al., 2020). Future research must expand empirical validation, integrate dynamic trust modeling, and assess trust frameworks across diverse operational settings to ensure HMT systems are both adaptable and practically viable.
3.4. Shared Mental Model (SMM)
Key Ideas. SMMs play a crucial role in aligning team understanding, facilitating effective communication, and enhancing HMT performance. SMMs can be categorized into task models, which focus on goals, required skills, and workload distribution, and role models, which define team members’ capabilities and responsibilities (Andrews et al., 2023). Empirical studies have shown that shared task and team mental models improve team dynamics and interactions, as demonstrated in flight-combat simulations (Mathieu et al., 2000). Research on HMT applications in USAR teams highlights the importance of shared cognition and communication efficiency, with natural language improving performance in Minecraft simulations (Demir et al., 2020). Industry 4.0 adaptation strategies further reinforce the role of design thinking and strategic alignment in team performance (de Paula et al., 2023). Proactive behaviors in robots, such as information sharing and joint task engagement, enhance team effectiveness under high cognitive loads (Edgar et al., 2023). Computational frameworks have also been proposed to enable artificial agents to develop and maintain SMMs, thereby improving coordination and team coherence (Scheutz et al., 2017).
Merits. Studies consistently demonstrate that well-structured SMMs enhance team coordination, communication, and performance across various HMT applications. Flight-combat simulations validate the influence of shared task and role models in improving team efficiency (Mathieu et al., 2000). The integration of proactive robot behaviors with SMMs in virtual reality simulations has shown promising results in mitigating cognitive overload and aligning human-robot goals (Edgar et al., 2023). In human-agent teams, planning has been shown to improve SMMs, enhancing coordination and reducing errors under high workloads (Stout et al., 1999). Furthermore, frameworks enabling artificial agents to maintain SMMs offer computational approaches for optimizing mixed human-agent teams (Scheutz et al., 2017). Studies on human-agent communication highlight the role of iterative cognition development and structured dialogue in improving decision-making and task execution (Schelble et al., 2022a).
Limitations. Despite strong theoretical and experimental support, research on SMMs in HMT faces notable limitations. Many studies rely on controlled simulations, such as Minecraft-based USAR tests (Demir et al., 2020) and virtual reality environments (Edgar et al., 2023), limiting their generalizability to real-world applications. Behavioral strategies for Industry 4.0 adaptation are based primarily on qualitative data, which lacks empirical validation (de Paula et al., 2023). The proposed computational frameworks for artificial agents do not fully capture the complexities of human cognition, which may hinder team coherence in real-world scenarios (Scheutz et al., 2017). In addition, human-agent team studies often fail to assess agents’ mental models directly, restricting deeper insights into HAT cognition and decision-making processes (Schelble et al., 2022a). To address these gaps, future research must integrate real-world testing, refine computational models for SMM adaptation, and explore scalable strategies for improving shared cognition in dynamic environments.
3.5. Team Situation Awareness
Key Ideas. Team Situation Awareness (TSA) is crucial for HMT performance, shaping communication, knowledge sharing, and coordination in dynamic environments. It aggregates individual situational awareness, where shared knowledge enhances stability, while dynamic settings require timely coordination (Salas et al., 2017). Measuring TSA remains challenging, as methods like SAGAT provide robust assessments but risk disrupting task dynamics (Endsley, 1995). UAV-based HMT studies show proactive information sharing improves TSA, though synthetic environments and language constraints limit applicability (Demir et al., 2019, 2017a). RPAS-based research highlights TSA’s role in performance, though machines struggle with non-textual data (McNeese et al., 2021c). Transparency models like the SAT framework enhance human-AI collaboration but introduce trade-offs in workload and decision latency (Chen, 2017). TSA models for collaborative driving, such as the SSA framework, improve trust alignment but assume hierarchical relationships, limiting real-world applicability (Kridalukmana, 2020).
Merits. TSA research offers valuable insights into how shared cognition and proactive communication enhance coordination in HMT. Transparency in AI decision-making improves human understanding, as demonstrated by the SAT model’s impact on trust and collaboration (Chen, 2017). Proactive information sharing strengthens TSA effectiveness in UAV- and RPAS-based HMTs, reinforcing structured communication’s role in team performance (Demir et al., 2019; McNeese et al., 2021b). Empirical findings highlight that TSA evolution in synthetic RPAS teams improves coordination over time (McNeese et al., 2021b). Collaborative driving research reduces automation over-reliance by aligning human trust with system capabilities (Kridalukmana, 2020). TSA research continues to refine decision-support frameworks, laying the foundation for scalable, real-time solutions.
Limitations. TSA research in HMT faces key limitations. Many studies rely on controlled simulations, such as UAV-based synthetic environments (Demir et al., 2019) and RPAS-based testing (McNeese et al., 2021b), limiting real-world applicability. Text-based machine agents struggle with multimodal interactions, reducing effectiveness in dynamic teams (McNeese et al., 2021c). TSA measurement methods like SAGAT must balance accuracy with operational flow, raising ecological validity concerns (Salas et al., 2017). Transparency models such as SAT enhance understanding but increase cognitive workload and decision latency, potentially impacting real-time performance (Chen, 2017). The SSA model assumes hierarchical team structures, oversimplifying automation trust dynamics where goal alignment and conflict resolution play critical roles (Kridalukmana, 2020). Future research should integrate real-time TSA assessment, improve machine adaptability, and ensure transparency mechanisms support rather than burden human decision-making.
3.6. Synergistic Human-AI Collaboration
Key Ideas. HMT enhances cyber malware detection by integrating human cognition with machine learning for pattern recognition. Expert knowledge is converted into visual ontologies and graph structures, enabling collaborative analysis of malware distribution networks (MDNs) (Cai et al., 2020). Human feedback refines machine learning algorithms, improving malware detection, mapping, and prediction while identifying critical vulnerabilities like hubs and bridges. This approach leverages human expertise and AI analytics to strengthen cyber defense strategies and adaptive threat mitigation.
Merits. The study demonstrates the benefits of combining human intelligence with AI to enhance malware detection and analysis. Visual ontologies and graph structures offer an intuitive framework for malware analysis, improving interpretability for both humans and machines (Cai et al., 2020). Human-in-the-loop feedback enables adaptive learning, allowing systems to refine detection capabilities over time. This integration supports proactive cybersecurity measures, highlighting HMT’s potential in complex threat environments.
Limitations. Despite its strengths, this approach faces scalability challenges due to reliance on high-quality human input, limiting automation and requiring expert involvement (Cai et al., 2020). Dependence on structured ontologies and graph-based models may hinder adaptability to evolving malware techniques. Data privacy concerns and computational costs further constrain implementation. Future research should reduce human dependency while maintaining accuracy, improve AI adaptability, and optimize computational efficiency for large-scale cybersecurity applications.
3.7. Team Communication & Coordination
Key Ideas. Effective communication is critical for collaboration in HMT, enhancing situational awareness, coordination, and adaptability (Endsley and Kaber, 1999). Transparent interactions foster trust and improves system performance (Lee and See, 2004), but existing strategies often fail to adapt to evolving technologies and user needs (Parasuraman et al., 2008). Research on search and rescue missions indicates that structured, on-demand explanations are as effective as frequent communication in improving situational awareness (Chiou et al., 2021). NLP-enhanced autonomous agents improve operational interactions but remain vulnerable to failures such as automation breakdowns and cyber-attacks (Demir et al., 2019). Studies on verbal interactions in HATs suggest that increased exchanges improve performance; however, these findings are primarily based on controlled simulations and need validation in real‐world settings (Demir et al., 2017b).
Merits. Research consistently highlights communication as a key factor in trust, situational awareness, and coordination. Optimized communication strategies enhance team efficiency, with moderate information exchange yielding the best performance (Demir et al., 2019). Machine learning-driven interfaces improve interaction quality and reduce cognitive burdens on operators (Moore, 2007). In high-stakes environments, selective communication strategies balance information delivery, preventing overload while maintaining awareness (Chiou et al., 2021). Advances in speech-based HMIs enable cooperative behavior through mutual intention recognition, addressing limitations in conventional systems (Moore, 2007).
Limitations. Many studies rely on simulations, limiting applicability to dynamic real-world environments (Chiou et al., 2021; Demir et al., 2017b). Research often assumes rigid hierarchical automation structures, failing to address adaptive interactions (Endsley and Kaber, 1999). Although NLP-enhanced systems improve interactions, they present usability challenges for untrained users—necessitating improved interface design and additional training (Demir et al., 2019). Speech-based HMIs demand precise parameter tuning, complicating large-scale deployment (Moore, 2007). Verbal behavior studies lack validation across diverse scenarios, necessitating empirical research on adaptive communication frameworks and scalable solutions for varying expertise levels.
3.8. Team Composition
Key Ideas. Human-autonomy teaming (HAT) research highlights challenges in communication, coordination, and team cognition. AI struggles with natural language processing and does not effectively leverage nonverbal cues, which humans rely on heavily (McNeese et al., 2021d). NeoCITIES experiments show AI-inclusive teams achieve higher TSA, while all-human teams exhibit superior cognition (McNeese et al., 2021d). Coordination effectiveness varies by environment, with synthetic teams showing performance fluctuations under shifting conditions (Demir et al., 2018). UAS task studies confirm all-human teams often outperform AI-integrated teams due to stronger coordination (McNeese et al., 2018). Team behaviors shift when humans are informed of AI roles, influencing social interactions but not necessarily improving performance (Walliser et al., 2019). AI’s impact depends on context, shaping team preferences and decision-making (Flathmann et al., 2023).
Merits. Research shows AI enhances TSA and structured coordination in HATs, particularly with reinforcement learning-based decision support (McNeese et al., 2021d). RPAS and CERTT-II studies confirm AI improves coordination in structured teams under stable conditions (Demir et al., 2018). Human perceptions of AI shape interaction strategies, reinforcing the importance of trust calibration in HMT (Walliser et al., 2019). AI’s adaptability in shared resource management influences decision-making (Flathmann et al., 2023), improving understanding of team composition, communication, and coordination in human-machine collaboration.
Limitations. AI struggles with naturalistic communication, lacking the ability to process implicit nonverbal cues (McNeese et al., 2021d). While AI improves TSA, it does not necessarily enhance cognition or performance in dynamic environments (McNeese et al., 2018). Many studies rely on controlled test beds, limiting real-world applicability (Demir et al., 2018). The impact of team-building interventions on HAT performance remains unclear, focusing more on behavior than quantifiable performance gains (Walliser et al., 2019). The absence of standardized metrics for assessing AI’s influence on shared decision-making complicates generalization across domains (Flathmann et al., 2023). Future research should integrate real-world validation, refine AI communication strategies to include nonverbal dynamics, and develop robust frameworks for long-term AI impact assessment in HMT.
3.9. Task Allocation
Key Ideas. Research on function allocation in HMT explores workload distribution, situational awareness, and responses to automation failures to enhance collaboration (Roth et al., 2019). While traditional frameworks guide allocation, they struggle with highly autonomous systems. Studies debunk full automation myths, emphasizing human roles in AI-assisted tasks such as content moderation and image labeling (Gray and Suri, 2019). This ghost work reveals how automation reshapes jobs rather than eliminating them, challenging conventional Human-Centered AI (HCAI) models. Task planning frameworks integrating human and robotic roles focus on balanced workload models and decision-making, optimizing hybrid workplace designs (Tsarouchi et al., 2017). These approaches automate task allocation while incorporating user-defined criteria.
Merits. Task allocation HMT research aims to optimize human-machine role distribution via workload balance and adaptability for efficiency (Roth et al., 2019). Recognizing human involvement in AI-assisted work reframes automation’s impact, showing AI transforms rather than replaces jobs (Gray and Suri, 2019). Task planning frameworks integrate human decision-making into automated processes, ensuring adaptability in dynamic settings (Tsarouchi et al., 2017). Multicriteria decision-making in task allocation enables scalable, user-driven hybrid workplace designs, boosting productivity while maintaining human oversight.
Limitations. Traditional function allocation models struggle to keep pace with autonomous systems, limiting long-term applicability (Roth et al., 2019). The persistence of human roles in automation underscores the need for transparency in AI decision-making and ethical considerations of ghost work (Gray and Suri, 2019). Existing task planning frameworks often lack scalability, requiring customization for different operational settings (Tsarouchi et al., 2017). Balancing human and machine contributions remains challenging, as rigid role allocation frameworks may reduce system flexibility. Future research should focus on adaptable, human-centric design principles to sustain function allocation in HMT environments.
3.10. AI & Ethics
Key Ideas. Ethical AI in HMT relies on responsibility, trustworthiness, transparency, and fairness to align with human values (Assaad and Boshuijzen-van Burken, 2023; Smith, 2019). While ethical compliance fosters trust, violations, especially regarding proportionality, undermine it (Textor et al., 2022). Trust in AI’s ethical decisions strengthens team cohesion and risk-taking, yet unethical behaviors disrupt coordination (Lopez et al., 2023). Research highlights the difficulty of trust repair in HATs, where even corrective actions like apologies fail to restore confidence (Schelble et al., ). Ethical frameworks such as Utilitarianism and Deontology guide AI decision-making but lack proven applicability in operational settings (Steen et al., 2023). While these frameworks define ethical AI interactions and transparency, limited empirical validation raises concerns about scalability and adaptability across HMT contexts (Smith, 2019; Assaad and Boshuijzen-van Burken, 2023).
Merits. Integrating ethical principles enhances AI accountability and ensures alignment with human values, reinforcing trust in HMT interactions (Assaad and Boshuijzen-van Burken, 2023). Ethical compliance sustains confidence in AI teammates, fostering collaboration and improving team dynamics (Textor et al., 2022; Lopez et al., 2023). Transparent, interpretable AI decision-making benefits from natural language processing, strengthening human-AI communication (Smith, 2019). Studies on MHC frameworks provide insights into balancing AI autonomy with human oversight in decision-making processes (Steen et al., 2023).
Limitations. Despite advances in ethical AI principles, translating them into scalable, validated strategies remains challenging across diverse HMT scenarios (Smith, 2019; Assaad and Boshuijzen-van Burken, 2023). Many studies rely on conceptual models and simulations, limiting real-world applicability (Steen et al., 2023; Textor et al., 2022). Ensuring explainability and mitigating bias remain challenging, as current techniques focus on post‐hoc explanations rather than proactive, fairness‐aware AI design. Bias in AI teammates can erode trust, particularly in high-stakes applications affecting safety and autonomy. Ethical trade-offs arise where AI autonomy for efficiency conflicts with human oversight, accountability, and fairness, especially in the military and healthcare. Trust repair remains difficult, as AI apologies fail to restore confidence in unethical AI teammates (Schelble et al., ). The lack of standardized evaluation metrics hampers benchmarking, raising concerns about adaptability in dynamic environments (Lopez et al., 2023). Future research must emphasize real-world validation, develop fairness-aware bias mitigation, and establish ethical frameworks that balance autonomy, accountability, and oversight in AI-driven HMT systems.
3.11. Discussions: Evolution of Research Focus in HMT
Fig. 6 shows research trends across ten HMT categories from 2017 to 2024, based on 93 studies. The distribution highlights a steady rise in activity across all areas, reflecting the growing interest in HMT systems.
Team Training consistently leads as the most researched area, emphasizing its central role in building effective HMT systems. Trust & Decision-Making is another prominent focus, highlighting its critical importance in fostering reliable collaboration between humans and machines. Other areas, such as Shared Mental Models, Team Autonomy, and Team Communication & Coordination, have shown steady growth, particularly in their contributions to improving interaction quality and alignment in team dynamics.
Figure 6. Frequency of ten categories of HMT research from 2017 to 2024.
Emerging areas like AI & Ethics and Synergistic Human-AI Collaboration have gained traction in recent years, addressing the increasing need for ethical considerations and user-centric design in HMT systems. The interconnectedness of these subsections is evident. Advancements in Team Training often support Trust & Decision-Making and Shared Mental Models, which influence Team Autonomy and Task Allocation. Likewise, ethical considerations in AI & Ethics impact Synergistic Human-AI Collaboration and Trust, highlighting the integrative nature of HMT research. These trends reflect evolving priorities and the synergy between HMT’s core areas, driving comprehensive system development. For reference, the contributions and limitations of each work are summarized in Table I of Appendix A (Empirical Studies to Promote Team Performance: Summary).
4. Evaluation Methodologies of Human-Machine Teaming Systems (HMTSs)
4.1. Evaluation Testbeds
Table 3 summarizes four evaluation methods commonly used for Human-Machine Teaming (HMT) systems: Statistics-based, Simulation Model-based, Emulation Testbed-based, and Real Testbed-based evaluations. Each method is described alongside examples and relevant references.
The Statistics-based Evaluation uses methods like Mixed Analysis of Variance (ANOVA) and Multivariate Analysis of Variance (MANOVA) to analyze data, focusing on performance measurement, trend analysis, error rates, and hypothesis testing (Chen and Barnes, 2014; Demir et al., 2020; Fotouhi et al., 2019; Javaid et al., 2023; McNeese et al., 2021c, d, 2018; Walliser et al., 2019; Murphy and Schreckenghost, 2013; Demir et al., 2021b; Damacharla et al., 2018; Gomez et al., 2019; Kashima et al., 2022; Yang et al., 2022; O’Neill et al., 2022; Berretta et al., 2023). This approach excels in control but is less realistic due to its reliance on mathematical models rather than real-world scenarios.
The Simulation Model-based Evaluation employs computational models to mimic real-world processes, allowing assessment of behavior and performance under various conditions (Demir et al., 2018; Cummings et al., 2010, 2011; McNeese et al., 2021c; Fotouhi et al., 2019; Javaid et al., 2023; Gay et al., 2019; Johnson et al., 2020; Lawless, 2021; Lawless et al., 2023; Liang et al., 2017; McDermott et al., 2005; Bersani et al., 2023; Flathmann et al., 2023; Nourbakhsh et al., 2005; Talamadupula et al., 2010; Wu et al., 2022; Murphy and Schreckenghost, 2013; Yang et al., 2022). Examples include testbeds like CERTT-II, NeoCITIES, self-driving car simulation platforms, and Minecraft. This method balances scalability and control but may lack the realism of physical environments. JADE-based frameworks like HACO provided development environments for simulating human-AI team interactions and testing teaming concepts (Dubey et al., 2020). The Emulation Testbed-based Evaluation creates controlled, realistic conditions for testing system performance and human-machine interactions. Examples include RPAS-STE, Malware Distribution Network (i.e., MDN) Graphs (Cai et al., 2020), RoboLeader (Chen and Barnes, 2014), Security Information and Event Management (SIEM) platforms (Gomez et al., 2019), Amazon Mechanical Turk (MTurk) (Kashima et al., 2022), CERTT-UAS-STE (i.e., Cognitive Engineering Research on Team Tasks - Unmanned Aerial System Synthetic Task Environment) (O’Neill et al., 2022), NASA-TLX, and COBALT (Johnson et al., 2021; Hart, 2006; Walliser et al., 2019; Knox and Stone, 2009; León et al., 2013; Nagy and Miller, 2021; Blaha et al., 2020). This approach combines realism with control, making it effective for detailed studies of HMT systems. The Real Testbed-based Evaluation focuses on real-world scenarios to examine human-AI interactions, such as Voice AI Agent and Aircraft with Embedded AI (Moore, 2007; Fotouhi et al., 2019; Lopez et al., 2023; Chen and Zimbra, 2010; Sheridan, 2016). While highly realistic and applicable, this method can be limited by scalability and logistical constraints. Table 3 outlines the methods, aiding in selecting suitable evaluation strategies for HMT systems based on research objectives and constraints. Detailed explanations of the cited works are available in the supplementary document (see Appendix B: Evaluation Methods for Human-Machine Teaming Systems).
Figure 7. Key attributes of HMT evaluation methods.
Comparative Analysis of the Four Evaluation Methods. Fig. 7 illustrates a comparative analysis of four evaluation methods used in Human-Machine Teaming (HMT) systems – Statistics-based, Simulation Model-based, Emulation Testbed-based, and Real Testbed-based – across five key criteria: realism, control, scalability, data quality, and applicability. Each evaluation method is represented as a distinct colored area, showcasing its strengths and trade-offs. Statistics-based evaluation scores are highly in control (5) due to their precise management of variables but have lower realism (2), reflecting its limited ability to replicate real-world scenarios. Simulation Model-based evaluation excels in scalability (5), making it suitable for diverse experimental conditions while maintaining moderate realism (3). Emulation Testbed-based evaluation effectively balances multiple criteria, achieving high data quality scores (5) and realism (4), allowing for controlled yet realistic assessments. Real Testbed-based evaluation provides maximum realism (5) and applicability (5) by operating in real-world settings, though it faces challenges in scalability (2) due to logistical constraints. Quantitative examples highlighted on the chart further illustrate these trends, emphasizing strengths such as high data quality in emulation-based methods and the unmatched realism of real testbeds. This visualization underscores the trade-offs inherent in each method, aiding researchers in selecting the most appropriate approach based on their evaluation objectives.
Table 3. Summary of HMT Evaluation Methods
Evaluation Method Description Examples with References
Statistics-based Evaluation Uses statistical methods like Mixed ANOVA and MANOVA for analyzing data, measuring performance, and testing hypotheses. Mixed ANOVA (Demir et al., 2020; McNeese et al., 2021c, d, 2018; Walliser et al., 2019), MANOVA (Demir et al., 2020; McNeese et al., 2021c; Demir et al., 2021b)
Simulation Model-based Evaluation Simulates real-world processes using computational models, assessing behavior and performance under varied conditions. CERTT-II Testbed (Demir et al., 2018; McNeese et al., 2018; Demir et al., 2017b; McNeese et al., 2021c), NeoCITIES (McNeese et al., 2021d; Schelble et al., 2020, 2022b, 2022a), ML-based Simulation (Bersani et al., 2023), Rocket League Platform (Flathmann et al., 2023), QN-MHP Testbed (Liu et al., 2006), Minecraft (Chiou et al., 2021; Demir et al., 2020), USAR Arenas (Nourbakhsh et al., 2005; Talamadupula et al., 2010)
Emulation Testbed-Based Evaluation Simulates realistic conditions using testbeds for controlled assessments of system performance and human-machine interactions. RPAS-STE (Johnson et al., 2021; Demir et al., 2019; Myers et al., 2019; McNeese et al., 2021a; Demir et al., 2021b), NASA-TLX (Hart, 2006; Gay et al., 2019), Strike Group Defender (Walliser et al., 2019), F-16 Fighter Aircraft (Mathieu et al., 2000), COBALT (Blaha et al., 2020)
Real Testbed-based Evaluation Tests HMT applications in real-world environments to study human-AI interactions under practical conditions. Voice AI Agent (Moore, 2007), Aircraft with Embedded AI (Lopez et al., 2023), AI and Opinion Mining (Chen and Zimbra, 2010)
4.2. Metrics
Table 4 presents an overview of the metrics used to evaluate HMT performance, categorized into three main types: human-centric, machine-centric, and team-centric metrics. Each category focuses on different aspects of HMT systems, providing complementary insights into their performance and dynamics.
Human-Centric Metrics emphasize evaluating human factors, including trust, collaboration, cognitive processes, and emotional responses. These metrics, such as perceived trust in AI, workload (e.g., NASA-TLX), and electrodermal activity (EDA), are crucial for understanding how humans interact with and perceive AI systems. They help identify factors influencing human engagement, workload, situational awareness, human reliability, and trust, which are critical for successful HMT operations (Chen and Barnes, 2014; Lopez et al., 2023; Walliser et al., 2019; Cummings et al., 2010, 2011; Ciechanowski et al., 2019; Gay et al., 2019; Johnson et al., 2020; Lawless et al., 2023; León et al., 2013; Chiou et al., 2021; McDermott et al., 2005; Nagy and Miller, 2021; Sheridan, 2016; Murphy and Schreckenghost, 2013; Edwards et al., 2023; Damacharla et al., 2018; Gomez et al., 2019; Kashima et al., 2022; Gebru et al., 2022; O’Neill et al., 2022). Error rates in AI decision-making and fault tolerance during system operation are often used as machine-centric metrics (Nagy and Miller, 2021). The ProtoEM method employs confusion matrices and the EM algorithm to estimate human uncertainty when integrating human and machine decisions in legal case matching tasks (Huang et al., 2024). Moreover, Knox and Stone (2009) evaluate the effectiveness of human trainers in shaping an agent’s behavior through reinforcement signals. These metrics primarily assess how human feedback influences the agent’s learning process, performance, and human-machine collaboration.
Machine-Centric Metrics assess the performance and reliability of machines or AI systems in HMT. Metrics such as response time, mission success rate, prediction accuracy, agent reliability, task allocation efficiency, transparency and explainability (i.e., how well AI explains its decisions), and task interpretation accuracy provide objective and reproducible measures of a machine’s capabilities. These metrics ensure that AI systems meet functional requirements and operate effectively in dynamic environments (Moore, 2007; Bersani et al., 2023; Cai et al., 2020; Cummings et al., 2010, 2011; Javaid et al., 2023; Liang et al., 2017; Patel et al., 2018; Beetz et al., 2017; Wu et al., 2022; Murphy and Schreckenghost, 2013; Chen and Barnes, 2014).
Team-Centric Metrics evaluate the collective performance and dynamics of human-machine teams. Metrics like situation awareness, team communication flow, shared mental models, human-agent collaboration efficiency, collaboration effectiveness, and creativity measure the synergy between human and AI agents, offering insights into collaboration, adaptability, and overall team efficiency. These metrics are essential for understanding how well the team functions as an integrated unit (Damacharla et al., 2018; Gay et al., 2019; Hauptman et al., 2023; Schelble et al., 2022a; Lawless, 2021; Stout et al., 1999; Hwang and Won, 2021; Murphy and Schreckenghost, 2013; Kashima et al., 2022).
Table 4 highlights the diverse range of metrics available for evaluating HMT systems, emphasizing their complementary roles in providing a comprehensive assessment. Combining these metrics can enhance the evaluation process, ensuring a balanced analysis of human, machine, and team performance. Detailed explanations of each work utilizing the metrics mentioned above are provided in the supplementary document (see Appendix C: Metrics for HMT Systems).
Comparative Analysis of Metrics for Evaluating HMT Systems. Fig. 8 comparatively analyzes the three categories of metrics used for evaluating HMT systems.
Figure 8. Key attributes of the three primary metrics for HMT systems.
Fig. 8 visualizes their performance across five key evaluation criteria: realism, subjectivity, complexity, scalability, and applicability. In realism, Team-Centric Metrics excel due to their focus on collaborative dynamics, while Human-Centric Metrics score high by capturing human behavioral nuances. Regarding subjectivity, Human-Centric Metrics rely heavily on subjective assessments, whereas Machine-Centric Metrics are largely objective, offering consistent and reproducible results. Complexity is another critical factor for both Human-Centric and Team-Centric metrics because they rely on sophisticated data collection and analysis methodologies. Scalability varies significantly across the categories; Machine-Centric metrics are the most scalable because of their quantitative nature. Human-Centric and Team-Centric metrics often face challenges when applied to diverse scenarios. Applicability is a strength for both Machine-Centric and Team-Centric metrics, as they are versatile across various HMT contexts. In contrast, Human-Centric metrics tend to be more specific to the context of human-AI interactions. This visualization underscores the complementary strengths and trade-offs of these metric categories, highlighting the importance of integrating them for a comprehensive evaluation of HMT systems.
4.3. Datasets
We highlight several notable trends and limitations in the datasets utilized for HMT research, as reviewed in this work. First, structured text datasets dominate (Cai et al., 2020; Fotouhi et al., 2019; Javaid et al., 2023; Gay et al., 2019; Johnson et al., 2020; Lawless, 2021; León et al., 2013; McDermott et al., 2005; Nagy and Miller, 2021), often capturing survey or interview-based insights into human-machine interactions.
Figure 9. Heatmap showing the distribution of dataset types and their availability in studies related to HMT systems.
However, many of these datasets remain unavailable or restricted, hindering reproducibility and progress in the field. Second, text datasets from sources like ENISA and NCSC (Mahaini et al., 2019) and audio datasets (Moore, 2007) are underrepresented. Text datasets are often secondary and collected from public repositories, while audio datasets are rarely used, limiting multi-modal interaction analysis. Third, numerical datasets, such as Lyapunov exponents (Demir et al., 2018), COBALT task data (Blaha et al., 2020), simulation data from RoboLeader (Chen and Barnes, 2014), crowdsourcing performance data (i.e., MTurk) (Kashima et al., 2022), physiological and psychological data (i.e., electroencephalography) (Ho et al., 2018), and human-automation interaction data (Cummings et al., 2010, 2011), are application-specific and cater to tasks requiring precise quantitative analysis. Speech and language processing datasets are vital for HMT systems, enabling the evaluation of human-machine communication and interaction quality (Berkol and Demirtaş, 2023). However, their restricted accessibility limits the replication and validation of results. Fourth, most datasets, including structured text and numerical data, are unavailable or partially accessible (Zhang et al., 2020). This lack of data availability challenges reproducibility and cross-study comparisons, restricting broader adoption of methodologies. Lastly, the reliance on structured text and numerical data reveals a lack of diverse datasets, such as video or sensor data, which could provide richer insights into complex human-machine interactions.
Making datasets publicly available fosters reproducibility and addresses these issues. Expanding to multi-modal data, such as video and audio, broadens research capabilities. Standardized formats and repositories tailored to HMT research could greatly enhance accessibility and advance the field.
To facilitate easy reference to the datasets used in each HMT study, we summarize their types and corresponding descriptions in Table 2 of Appendix B (Metrics for HMT Systems).
Dataset Availability and Distribution in HMT Research. Fig. 9 presents dataset distribution and availability across HMT studies, categorizing Text, Audio, Structured Text, and Numerical Data as Available, Partially Available, or Unavailable. Structured text is the most used but often unavailable, limiting reproducibility and adoption. Numerical and text datasets contribute but remain less accessible. The ”Low” and ”High” labels on the color bar indicate dataset frequency, with darker shades representing higher occurrences. This visualization underscores dataset usage trends, emphasizing the need for improved accessibility and shared repositories.
Table 4. Metrics for Evaluating Human-Machine Teaming (HMT) Performance
Metric Type Description References
Human-Centric Metrics Metrics focusing on human factors in HMT performance evaluation, including trust, collaboration, and cognitive measures. (Lopez et al., 2023; Walliser et al., 2019; Damacharla et al., 2018; Cummings et al., 2008; Ciechanowski et al., 2019; Gay et al., 2019; Knox and Stone, 2009; Liu et al., 2006; Schelble et al., 2022a; Fallon and Blaha, 2018; Sheridan, 2016; Chiou et al., 2021)
Machine-Centric Metrics Metrics assessing machine or AI system performance, such as response time, mission success rate, and accuracy in classification tasks. (Moore, 2007; Bersani et al., 2023; Fallon and Blaha, 2018; Cai et al., 2020; Chiou et al., 2021; Damacharla et al., 2018; Yildirim, 2022; Knox and Stone, 2009; Blaha et al., 2020; Liang et al., 2017; Patel et al., 2018; Lee et al., 2015; Wu et al., 2022; Beetz et al., 2017)
Team-Centric Metrics Metrics evaluating the performance and dynamics of human-machine teams, including situation awareness, communication flow, and shared mental models. (Damacharla et al., 2018; Hauptman et al., 2023; Demir et al., 2017b; Schelble et al., 2022a; Stout et al., 1999; Gay et al., 2019; Hwang and Won, 2021; McNeese et al., 2021c; National Academies of Sciences, Engineering, and Medicine and others, 2021)
5. Key Theoretical Approaches for Human-Machine Teaming
Through an extensive review of HMT systems, we have identified three widely utilized primary theories: Human-in-the-Loop Reinforcement Learning (HRL), Instance-Based Learning Theory (IBLT), and Interdependence Theory (IT).
5.1. Human-in-the-Loop Reinforcement Learning (HRL)
Key Ideas. AI alone is often less effective than human expertise in solving complex problems (Wu et al., 2022). To address this, Human-in-the-Loop Reinforcement Learning (HRL) integrates human knowledge into AI decision systems to enhance prediction accuracy at minimal cost. HRL research addresses (1) improving data processing by selecting relevant samples, (2) enhancing model performance via human-provided high-dimensional knowledge, and (3) designing efficient systems to incorporate human input without compromising performance or robustness.
HRL Approaches in HMT Systems. Liang et al. (2017) proposed an HRL framework for vehicle decision-making, integrating human-vehicle interactions with auxiliary driving equipment. They introduced a hybrid Markov decision process model leveraging human intuition. León et al. (2013) developed a reinforcement learning framework combining human demonstrations and feedback, enabling non-experts to train robots. When applied to a “pick and place” task, it improved learning efficiency, robustness to noisy data, and task performance. Knox and Stone (2009) introduced the Tamer framework, using human reinforcement to train agents without environmental rewards, reducing sample complexity and enabling rapid policy learning. While Tamer agents excel initially, autonomous agents outperform them with extensive training. A hybrid approach balancing human reinforcement and environmental rewards enhances efficiency and long-term performance in delayed-feedback scenarios.
Limitations. These HRL-based studies show promise in integrating human knowledge into AI systems but face challenges like scalability across diverse tasks, biases from human input, and limited validation in dynamic real-world environments. Future research should prioritize scalable frameworks, reducing biases, enhancing adaptability, and validating performance in realistic settings to ensure trustworthiness in HMT systems.
5.2. Instance-based Learning Theory (IBLT)
Key Ideas. Instance-Based Learning Theory (IBLT) (Gonzalez et al., 2003) explains learning in dynamic decision-making (DDM) through accumulating, recognizing, and refining instances. Knowledge is stored as Stored Decision Units (SDUs), capturing context, actions, and expected outcomes. Decision-making relies on recognition-based retrieval, where memory retrieves SDUs based on situational similarity. Learning adapts through adaptive strategies, alternating between heuristics for unfamiliar situations and instance-based approaches for familiar ones. Necessity-based choice balances exploring alternatives and committing to the best-known option based on constraints like time or preference. Feedback updates refine SDU utilities by linking actions to observed outcomes, enhancing decision accuracy over time.
IBLT Approaches in HMT Research. Gonzalez et al. (2022) tackled static cyber defense limitations, proposing an adaptive IBLT framework for dynamic deception using signaling, masking, and decaying. Integrating game theory and machine learning, this model refined cognitive strategies but lacked empirical validation for safety and trust. Instance-Based Learning (IBL) generates predictions from individual instances without abstraction (Aha et al., 1991). Gupta et al. (2023) extended the TSM-CI framework, identifying IBLT as a suitable AI architecture. Dong et al. (2022) introduced the Dispatching Multitasking Trust Paradigm (DMTP) to enhance dispatcher trust in Fully Automated Operation (FAO) metro systems, though its IBLT-based trust model lacked validation. Yang et al. (2023) explored human-machine communication via Machine Theory of Mind, proposing three MToM approaches: modeling human inference, simulating AI-like behavior, and integrating domain-specific actions. Their IBLT scenario demonstrated effective collaboration. Blaha et al. (2020) developed an IBLT-based computational model of human decision-making, showing how trust calibration evolves via task experience.
Limitations. While IBLT offers valuable mechanisms for dynamic decision-making and trust calibration, its application in developing trustworthy HMT systems faces notable limitations. Existing studies often lack empirical validation, particularly in real-world dynamic environments, which undermines their scalability and practical applicability. Further, the reliance on predefined instances limits adaptability in unpredictable scenarios, where novel situations demand innovative solutions. Future research should incorporate empirical evaluations, refining feedback mechanisms, and enhancing adaptability by integrating IBLT with advanced cognitive architectures to address the complexities of HMT.
5.3. Interdependence Theory (IT)
Key Ideas. Independence theory (IT) in HMT emphasizes balancing human and machine autonomy to optimize collaboration by delineating decision-making and task boundaries, leveraging strengths, and minimizing conflicts (Sheridan, 2016). Combined with interdependence theory, it identifies tasks best suited for humans or machines, fostering efficiency and synergy for enhanced team performance beyond individual capabilities.
IT Approach in HMT Research. Johnson et al. (2020) emphasized interdependence analysis to enhance human-autonomy teaming, addressing the limitations of technology-centric tools and promoting system adaptability. Lawless et al. (2023) criticized traditional team science for treating individuals as independent, proposing state dependency to resolve validation and replication crises by examining interconnected team dynamics. Nagy and Miller (2021) highlighted the challenge of diagnosing ML algorithms due to limited code transparency, advocating for human interventions via interdependence analysis. Lawless (2021) argued interdependence improves team performance, countering the tendency to oversimplify data. However, Lawless (2020) noted the lack of a formal mathematical foundation for designing and operating Autonomous HMTs (A-HMTs), calling for mathematical models to enhance development and reliability.
Figure 10. Key benefits of HRL, IBLT, and IT for HMT performance.
Limitations. Existing HMT studies using independence theory rely on static task boundaries and insufficiently integrate dynamic interdependence in human-machine interactions (Sheridan, 2016; Lawless et al., 2023). They often overlook adaptive mechanisms for evolving team dynamics, ML safety concerns, and a mathematical foundation for interdependence modeling (Nagy and Miller, 2021; Lawless, 2020). Future research should develop formalized interdependence models to enhance adaptability, safety, and performance in A-HMTs, incorporating real-time feedback to adjust task allocations and roles based on team and context.
Comparative Analysis of HRL, IBLT, and IT Frameworks. Fig. 10 compares the strengths of HRL, IBLT, and IT across six key factors. HRL excels in decision-making efficiency, feedback integration, and human-machine interaction, making it ideal for real-time systems like autonomous vehicles or disaster response. IBLT offers superior adaptability by leveraging prior cases, making it suitable for healthcare diagnostics and training. IT is strongest in task allocation and overall performance, highlighting its relevance in collaborative applications like project management or organizational planning. HRL emphasizes real-time adaptability with human oversight, IBLT leverages historical knowledge, and IT focuses on interdependencies and coordination. HRL suits environments requiring continuous feedback, while IBLT is preferable for case-based reasoning. IT thrives in collaborative settings where role optimization is key. Each framework has distinct advantages, suggesting domain-specific applications with potential synergies for comprehensive solutions.
6. Concepts, Challenges, and Applications of Human-Machine Teaming
6.1. Human and Unmanned Vehicle Teaming
Concept of UAV Teaming. Unmanned Aerial Vehicle (UAV) teaming involves collaborative operations between UAVs or between UAVs and human operators to achieve shared objectives (Fotouhi et al., 2019). This concept leverages UAVs’ strengths, such as agility, autonomy, and scalability, to perform complex tasks efficiently. Effective UAV teaming requires seamless integration of communication, coordination, and control to adapt to dynamic and uncertain environments (Javaid et al., 2023). McDermott et al. (2005) and Cummings et al. (2011) emphasized the importance of optimizing human-automation interaction for unmanned vehicle operations, highlighting how operator proximity and UV type influence mission efficiency and how decentralized planning can enhance effectiveness in dynamic scenarios.
Challenges in UAV Teaming Systems. Key challenges in UAV teaming include ensuring reliable communication, balancing workload between humans and automation, and maintaining coordination in dynamic environments. Cummings et al. (2010) highlighted that high automation replanning rates increase operator workload, while Calhoun et al. (2018) proposed adaptive interfaces to improve decision-making, though both faced real-world validation issues. Additional challenges include scalability, adapting to unpredictable conditions, and resilience against cyber threats (Gay et al., 2019).
UAV Teaming Applications and Their Limitations. UAV teaming has been explored as follows:
  • Military Operations: UAV teaming supports surveillance, reconnaissance, and tactical missions. Gay et al. (2019) studied cyberattack detection in unmanned ground vehicles (UGVs), introducing “Suspicion Theory” to structure operator suspicion responses. While promising, this approach needs further adaptation to complex military environments.
  • Disaster Response: UAV teams enhance situational awareness and coordination in search-and-rescue missions. However, scalability and challenges in real-time coordination limit their broader applicability (McDermott et al., 2005).
  • Logistics and Supply Chain: UAVs are employed for package delivery and supply chain optimization. Yet, regulatory constraints, battery limitations, and collision avoidance remain significant hurdles (Cummings et al., 2011).
  • Automation Workload Balance: Research has proposed strategies to balance automation rates and operator workload in UAV networks (Cummings et al., 2010; Calhoun et al., 2018). However, these strategies require empirical validation in diverse, dynamic contexts.
Despite advancements, reliance on simulations that miss real-world complexities remains a key limitation. Limited focus on communication or automation rates underscores the need for integrated, scalable UAV teaming systems.
6.2. Human-Chatbot Teaming
Concept of Human-Chatbot Teaming (HCT). Human-Chatbot Teaming (HCT) refers to collaborative interactions where humans and chatbots work together to achieve shared objectives. This concept leverages chatbots’ strengths, such as accessibility, consistency, and rapid information processing, to enhance team productivity and decision-making (Chen et al., 2023; Ho et al., 2018). Effective HCT relies on integrating conversational interfaces with contextual understanding and adaptability to foster seamless collaboration in diverse environments (Mygland et al., 2021; Al-Shafei, 2024).
Challenges in HCT Systems. HCT systems face significant challenges, including achieving natural and context-aware conversations, managing user trust, and ensuring equitable workload distribution (Ho et al., 2018; Al-Shafei, 2024). Developing chatbots that can adapt to individual user preferences and maintain engagement in long-term interactions is particularly difficult (Mygland et al., 2021; Chen et al., 2023). Furthermore, technical challenges like integrating with diverse platforms, addressing ethical concerns such as bias, and ensuring data privacy and security add to the complexity of designing robust HCT systems (Al-Shafei, 2024). These challenges are amplified in dynamic and high-stakes scenarios where accuracy and reliability are critical (Ho et al., 2018; Chen et al., 2023).
HCT Applications and Their Limitations. HCT has been explored across various domains, with notable studies highlighting the following applications and limitations:
  • Emotional and Collaborative Responses: Ciechanowski et al. (2019) investigated human-chatbot interactions, revealing that simpler chatbots elicited fewer emotional responses, with users favoring less human-like features. This suggests that chatbot mimicry of human traits may not be essential for effective collaboration.
  • Task Performance Enhancement: Sowa and Przegalinska (2020) showed that chat-based virtual assistants enhance productivity in business simulations, highlighting chatbots’ potential to improve team outcomes through assistance.
  • Security Task Collaboration: Nagy and Miller (2021) developed a framework to improve vulnerability detection in software, finding that teams with virtual assistants outperformed standalone tools in error identification.
  • Creativity and Team Perception: Hwang and Won (2021) explored creativity in brainstorming, showing that conversational styles and perceived identities affect task performance, team dynamics, and creative self-efficacy.
Despite these advancements, common limitations persist. Many studies rely on controlled environments and simplified tasks, reducing real-world applicability. The lack of face-to-face communication in some experiments further limits generalizability. Besides that, framework-specific constraints and limited tool integration hinder broader adoption and scalability, emphasizing the need for more adaptable and comprehensive HCT systems.
6.3. Human-Robot Teaming
Concept of HRT. Human-Robot Teaming (HRT) refers to the collaboration between humans and robots to achieve shared goals by leveraging their complementary strengths (Natarajan et al., 2023; Ciocirlan et al., 2019). Effective HRT relies on communication, coordination, and trust to ensure seamless interaction and optimize performance in dynamic environments (Natarajan et al., 2023; Ciocirlan et al., 2019).
Designing effective HRT systems involves key elements: understanding agent abilities, taskwork, metrics, and peer-to-peer interactions. Agent abilities enable effective task execution, taskwork focuses on specific activities, metrics evaluate performance, and peer-to-peer interactions ensure communication and coordination. Balancing these factors is essential for collaboration and optimal outcomes (Mingyue Ma et al., 2018).
Challenges in HRT Systems. HRT faces challenges in developing robust communication, fostering trust, and adapting to dynamic tasks. Robots must predict human actions, and algorithms must handle evolving conditions. Trust is essential for human confidence in robot capabilities (Natarajan et al., 2023).
HRT Applications and Their Limitations. HRT applications cover diverse domains, including:
  • Humanoid Robots: Hoffman and Breazeal (2004) introduced Robonaut, a humanoid robot designed for space missions. It utilized the Joint Intention Theory for collaborative tasks but focused on theoretical aspects with limited empirical testing. Reliance on complex social cues may restrict applicability in non-social settings.
  • Simulation Testbed for HRT Systems: Nourbakhsh et al. (2005) and NIST developed a testbed for search-and-rescue operations using the multi-agent system Retsina. While effective for training, challenges included spatiotemporal coordination and complex control interfaces, hindering multi-robot management.
  • Planning Systems for HRT: Talamadupula et al. (2010) proposed planning systems for dynamic environments, integrating counterfactual reasoning and conditional goals. However, scalability, reliance on accurate environmental models, and robust communication protocols remain hurdles.
  • Team Coordination: Gombolay et al. (2015) explored human-robot collaboration with the Willow Garage PR2 platform. While prioritizing human preferences improved perception, it reduced overall task efficiency, highlighting trade-offs between satisfaction and performance.
6.4. Human-AI Teaming-based Social Media
Concept of HAT-based Social Media. HAT (Human-AI Teaming)-based Social Media refers to the collaborative integration of AI technologies and human decision-making processes to enhance the efficiency, accuracy, and ethical standards of social media systems. This concept leverages AI’s ability to process vast amounts of data, provide contextual insights, and support human judgment while maintaining critical oversight and ethical considerations (brundage2019toward; Hagendorff, 2023). Effective HAT-based social media systems aim to address challenges such as misinformation, content moderation, and user engagement by combining AI’s computational strengths with human creativity and ethical sensibility (Tufchi et al., 2023; Balayogi et al., 2025).
Challenges in HAT-based Social Media. HAT-based social media systems face numerous challenges, including balancing automation and human oversight, ensuring transparency in AI decision-making, and addressing biases in AI models. The complexity of social media data, such as informal language, cultural nuances, and real-time dynamics, complicates the development of robust AI systems. Besides that, ethical concerns like user privacy, algorithmic fairness, and the potential misuse of AI for malicious purposes amplify the difficulty of creating trustworthy HAT systems. Regulatory and governance challenges also emerge, requiring international cooperation and adaptable policies to ensure compliance with ethical standards and data protection laws.
HAT-based Social Media Applications and Their Limitations. Applications of HAT-based social media span a wide range of domains, with notable studies highlighting both their potential and limitations:
  • Fake News Detection: A Hybrid Multi-thread Metaheuristic method has been introduced for detecting fake news, demonstrating competitive performance on COVID-19 and Syrian war datasets (Yildirim, 2022). However, hardware constraints and limited dataset applicability present significant challenges.
  • Opinion Analysis: AI frameworks integrating sentiment analysis and stock modeling have been developed to assess user opinions on world events (Chen and Zimbra, 2010). Despite their innovative approach, external variables and unreliable sentiment extraction limit their accuracy.
  • Health Issue Detection: Ethical AI-integrated social media has been explored to detect health-related issues, ensuring fairness and privacy (Liu et al., 2023). However, findings remain constrained by small sample sizes and qualitative data collection.
  • Cultural Ecosystem Services (CES): AI-driven analysis of Flickr images has been used to assess CESs and identify key hotspots (Egarter Vigl et al., 2021). Broader applicability is limited by geographic constraints and ethical concerns surrounding data use.
  • Crisis Response: AI-based disaster-related social media analysis has been applied to enhance crisis response, proposing methods for improving detection and intervention (Imran et al., 2020). Challenges persist in detecting minor disasters and mitigating risks posed by manipulated images.
  • Social Media Marketing for SMEs: AI-driven social media marketing has been studied for its impact on customer retention and business growth (Basri, 2020). However, outdated analytics and vague survey methods limit the insights gained.
  • Regulation of AI Content: Regulatory solutions for AI-generated social media content have been proposed, including traceability logs for private data to enhance accountability (Lewis and Moorkens, 2020). However, governance receptivity and legal complexities continue to hinder implementation.
These applications demonstrate the potential of HAT-based social media systems to address critical issues, but limitations such as scalability, ethical challenges, and data biases underscore the need for continued research and development to achieve reliable and equitable outcomes.
6.5. HMT-based Smart Computing
Concept of HAuT-based Smart Computing. Human-Autonomy Teaming (HAuT)-based Smart Computing refers to the integration of autonomous systems and human decision-making to address complex computational tasks (Lyons et al., 2021). This concept leverages the strengths of autonomy, such as efficiency, scalability, and data-driven insights, while maintaining critical human oversight for ethical, adaptive, and context-aware decision-making (O’Neill et al., 2022). By combining human creativity with machine precision, HAuT-based smart computing aims to enhance productivity, reliability, and situational awareness across diverse domains like manufacturing, IoT, and swarm robotics (Abbass and Hunjet, 2021).
Challenges in HAuT-based Smart Computing. HAuT-based smart computing systems face significant challenges in scalability, adaptability, and maintaining mutual trust between humans and autonomous systems. Ensuring real-time data processing and decision-making in dynamic and unpredictable environments is particularly demanding (Abbass and Hunjet, 2021; Haindl et al., 2022a). The computational demands of relational machine learning and the intricacies of monitoring and validating ethical policies further complicate the implementation of robust systems (Haindl et al., 2022a, b). Moreover, the lack of adaptability to rapidly evolving technologies and maintaining effective communication in complex scenarios remain pressing issues (Haindl et al., 2022b; Oh et al., 2024).
HAuT-based Smart Computing Applications and Their Limitations. Applications of HAuT-based smart computing demonstrate significant potential but also reveal inherent limitations:
  • Human-Swarm Teaming: Abbass and Hunjet (2021) introduced intelligent shepherding agents to manage human-swarm teams, emphasizing contextual awareness for autonomous decision-making. While promising, the architecture requires validation using real data and precise metrics, especially in dynamic and unpredictable environments.
  • Smart Manufacturing: Haindl et al. (2022a) proposed a reference architecture for AI-enabled smart manufacturing, focusing on enhancing mutual trust and situational awareness. However, scalability for real-time data processing and the complexity of implementing and validating ethical policies remain challenges.
  • Software Platforms for HAuT: Haindl et al. (2022b) identified software platform characteristics critical for effective human-autonomy teaming, such as explicability and trustworthiness. However, scalability and adaptability issues hinder deployment in rapidly evolving manufacturing technologies.
  • IoT Environments: Oh et al. (2024) developed a rule-based adaptive system for improved human-machine collaboration in IoT setups, validated in simulated environments. However, maintaining effective communication and scalability in real-world scenarios remains a significant limitation.
These applications demonstrate the potential of HAuT-based smart computing across fields. However, ensuring scalability, adaptability, and ethical integrity is essential for robust and trustworthy implementations.
6.6. HMT-based Natural Language and Information Retrieval
Concept of HMT-based NLP and IR. HMT-based NLP and Information Retrieval (IR) integrate human expertise with machine intelligence for complex language understanding, extraction, and retrieval (Berkol and Demirtaş, 2023). By combining human decision-making and contextual insights with machine-driven data processing, this approach enhances accuracy, efficiency, and adaptability in applications such as taxonomy development, rescue missions, and legal analysis (Wan et al., 2022; Safdar et al., 2024; Mahaini et al., 2019; Beetz et al., 2017; Huang et al., 2024).
Challenges in HMT-based NLP and IR. HMT-based NLP and IR face challenges like scalability, adaptability to dynamic data, and balancing human input with machine automation. Manual data collection and refinement limit scalability (Mahaini et al., 2019), while unpredictable language expressions pose obstacles in real-world scenarios (Beetz et al., 2017). Ensuring robustness in diverse applications and integrating these systems into workflows adds complexity (Dubey et al., 2020; Edwards et al., 2023). Capturing tacit knowledge and addressing decision-making uncertainties remain critical hurdles (Huang et al., 2024).
HMT-based NLP and IR Applications and Their Limitations. Applications of HMT-based NLP and IR demonstrate significant potential but also reveal notable limitations:
  • Cybersecurity Taxonomy Development: Mahaini et al. (2019) used HMT to streamline cybersecurity taxonomy creation by integrating NLP and IR tools for term extraction, guided by human decision-making. However, reliance on manual processes limits scalability and adaptability to new domains.
  • Natural Language Command Interpretation in Rescue Missions: Beetz et al. (2017) developed NLP systems to help robots interpret human instructions in simulated rescue scenarios. While effective in controlled settings, the system’s performance is limited by unpredictable real-world elements and diverse language expressions.
  • Framework for Human-AI Collaboration: Dubey et al. (2020) extended the Java Agent Development Framework (JADE) to improve human-AI interactions through a structured taxonomy and graphical interfaces, reducing development time. However, scalability and robustness in diverse contexts remain key challenges.
  • Evidence Gap Map (EGM) Creation: Edwards et al. (2023) integrated AI, specifically a BERT-based model, to automate the initial screening in evidence synthesis processes, significantly reducing human workload. Yet, the approach faces challenges with larger datasets, dependency on high-quality training data, and workflow integration.
  • Legal Case Matching: Huang et al. (2024) proposed the Co-Matching framework, integrating human tacit knowledge and machine analysis to enhance legal document matching accuracy. Articulating tacit knowledge and estimating human decision uncertainty are key limitations, especially with limited behavioral data.
These applications highlight the promise of HMT-based NLP and IR in addressing complex challenges across various domains. However, scalability, adaptability, and the integration of human expertise into machine-readable processes remain critical areas for future research and development.
7. Conclusions & Future Work
7.1. Key Findings
This survey provides a comprehensive examination of HMT by integrating insights from computational, cognitive, and social sciences. Our findings highlight several critical advancements and challenges shaping the field.
Advancements in HMT. HMT research has significantly evolved across domains such as defense, healthcare, industrial automation, and autonomous systems. Advances in team training, shared mental models (SMMs), and trust calibration have contributed to improved team performance, decision-making, and adaptability. Studies demonstrate that AI-driven collaboration frameworks, such as human-in-the-loop reinforcement learning (HRL) and instance-based learning theory (IBLT), enhance adaptive teaming mechanisms and efficiency. Moreover, the development of explainable AI (XAI) techniques has strengthened human trust in automated decision-making.
Challenges in Trust, Transparency, and Ethical Considerations. Trust remains a key issue in HMT, particularly in explainability, reliability, and ethical AI decision-making. Research shows that trust calibration between human and AI teammates is dynamic and difficult to restore once compromised. Transparency models like situation awareness-based agent transparency (SAT) improve trust but increase cognitive workload. In addition, ethical frameworks guiding AI behavior remain largely theoretical, with limited empirical validation in real-world HMT applications.
Task Allocation and Decision-Making. Effective task allocation in HMT depends on balancing automation with human oversight. While reinforcement learning and multi-criteria decision-making enable scalable human-AI collaboration, challenges remain in role adaptability and real-time adjustments. Research shows that team composition, especially in mixed human-agent teams, impacts task performance, coordination, and cognitive alignment.
Evaluation Frameworks and Benchmarking. The need for standardized evaluation methodologies in HMT remains a critical challenge. While existing testbeds and simulations provide controlled insights into team dynamics, real-world applicability is limited. Studies leveraging NASA-TLX, CERTT-II, and NeoCITIES testbeds highlight gaps in scalability and generalizability. Moreover, dataset availability remains an obstacle, as structured text and numerical datasets dominate, restricting access to multi-modal and real-world interaction data.
7.2. Future Research Directions
Based on the identified challenges, future research in HMT should focus on the following key areas:
Enhancing Explainability and Trust Calibration. Future work should explore advanced XAI mechanisms that integrate transparency with adaptive learning models to improve user trust. Developing real-time trust assessment methods will enable HMT systems to dynamically adjust interactions based on evolving human-agent relationships.
Human-Centered AI Design. Research should emphasize user-centered design principles to improve HMT effectiveness by ensuring AI teammates align with human cognitive models, decision-making patterns, and workload considerations. Future studies should explore adaptive interfaces that enhance usability and foster intuitive collaboration.
Scalability and Real-World Deployment. Transitioning HMT systems from controlled settings to real-world applications remains challenging. Future research should explore scalable architectures that balance performance and adaptability while ensuring robust human-agent interaction within domain-specific constraints.
Standardized Evaluation Metrics and Testbeds. Establishing benchmarking standards for HMT systems is essential to ensure comparability across studies. Developing large-scale empirical testbeds that integrate real-world HMT scenarios will provide more robust performance evaluations. Future research should also address the reproducibility crisis in HMT by improving dataset accessibility and adopting open-source benchmarking frameworks.
Ethical and Regulatory Considerations. As HMT expands into high-stakes environments, future research must tackle ethical AI governance, bias mitigation, and accountability in autonomous decisions. Developing policy frameworks that align AI autonomy with human oversight is crucial for responsible HMT deployment.
Multi-Modal and Cross-Domain Applications. Future studies should move beyond text-based interactions to multimodal AI systems, including visual, auditory, and sensor-based communication. Expanding HMT to disaster response, cybersecurity, and human-robot teaming will improve adaptability and scalability.
Real-World Validation of HMT Theories. While theoretical models such as HRL, IBLT, and Interdependence Theory (IT) offer valuable insights, their real-world applicability remains largely unexplored. Future work should focus on validating these frameworks through longitudinal studies and large-scale empirical deployments.
By addressing these directions, HMT research can advance toward more resilient, ethical, and scalable human-machine collaboration models. Continued interdisciplinary research, industry-academic partnerships, and policy-driven AI governance will be essential in shaping the next generation of HMT systems.
Acknowledgment
This research is partially supported by the U.S. Army Research Office (ARO) Award (W911NF-24-2-0241) and the National Science Foundation (NSF) Secure and Trustworthy Cyberspace (SaTC) Award (2330940).